OpenStudy (anonymous):

Solve for t. √t+9=16

OpenStudy (experimentx):

49

OpenStudy (lgbasallote):

if you want the explanation...it's isolate t first... sqrt t +9 -9 = 16 -9 sqrt = 7 square both sides.. t = 7^2 t= 49